OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 3/2x + 4 and passes through the point (−6, 3).

OpenStudy (anonymous):

Gradient for the perpendicular line =-1/m = -2/3 Using y-y1=m(x-x1) or y=mx+b, find the equation of the perpendicular. Your m would be -2/3 and your (x1, y1) would be (-6, 3) Equation of perpendicular: 3y=-2x+21

OpenStudy (anonymous):

Don't forget to fan :$

OpenStudy (anonymous):

thats not on of the answers tho.

OpenStudy (anonymous):

y=-2x/3+7

OpenStudy (anonymous):

Does that work?

OpenStudy (anonymous):

y = 3/2x − 1 y =3/2 x + 3 y = −2/3x − 1 y = −2/3x + 3

OpenStudy (anonymous):

y = −2/3x − 1

OpenStudy (anonymous):

My bad, I made a negative sign error. Sorry.

OpenStudy (anonymous):

its okay. thank you!

OpenStudy (anonymous):

Fan :P

OpenStudy (anonymous):

I am.